OpenStudy (anonymous):

Use the addition formulae to find an expression for \[\cos^2(A+B) + \sin^2(A+B). Verify~that~your ~expression~ reduces \to 1.\]

OpenStudy (anonymous):

wow, according to the instructions, this could get long....

OpenStudy (anonymous):

Yes, it can... I got up to a step where cosA^2AcosB^2+sinA^2SinB^2+sinA^2cosB^2+cosA^2SinB^2.... But I have no clue if I'm correct...!

OpenStudy (anonymous):

cos^2A(sin^2B + cos^2B) + sin^2B(cos^2a + sin^2a)... I am sure you will get something like this after taking some terms common. #I would have left this question, the answer is obvious and the solution according to the instructions is pretty boring,

OpenStudy (anonymous):

I'd still like a run through to see if someone gets that answer, and then how to get 1 from there...

OpenStudy (anonymous):

ok so lets see if you're on the right track... cos(A+B) = cosAcosB - sinAsinB sin(A+B) = 2sinAcosB so cos^2(A+B) + sin^2(A+B) = (cosAcosB - sinAsinB)^2 + (2sinAsinB)^2 still with me?

OpenStudy (anonymous):

Yes, I am. Sorry, my internet is going on and off!

OpenStudy (anonymous):

pay no attention to my previous post. there was an error in the last statement. using sum rule for cosine and sine we get... (cosAcosB - sinAsinB)^2 + (2sinAcosB)^2 so FOILing this out we get: (cosAcosB)^2 - 2*(cosAcosB)(sinAsinB) + (sinAsinB)^2 +4sin^2Acos^2B

OpenStudy (anonymous):

(cosAcosB)^2 - 2*(cosAcosB)(sinAsinB) + (sinAsinB)^2 +4sin^2Acos^2B cos^2A*cos^2B - 2*cosA*cosB*sinA*sinB + sin^2A*sin^2B +4*sin^2A*cos^2B

OpenStudy (anonymous):

cos^2A*cos^2B - 2*cosA*cosB*sinA*sinB + sin^2A*sin^2B +4*sin^2A*cos^2B cos^2A*cos^2B - 2*cosA*cosB*sinA*sinB + (sin^2A*sin^2B +4*sin^2A*cos^2B) cos^2A*cos^2B - 2*cosA*cosB*sinA*sinB + sin^2A(sin^2B +4*cos^2B) this is where I am so far. gimme a sec.... lemme do it on paper...

OpenStudy (anonymous):

Ok. Sure. No problem.

OpenStudy (anonymous):

ok, no wonder I wasn't getting anywhere.... sin(A+B) = sinAcosB + cosAsinB sorry 'bout that. continuing...

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

so, \[\cos ^{2}(A+B)=(cosAcosB - sinAsinB)^{2}\] and \[\sin ^{2}(A+B) = (sinAcosB + cosAsinB)^{2}\]

OpenStudy (anonymous):

FOILing out the right side of both of these and adding them together will give you \[(cosAcosB)^{2}+(sinAcosB)^{2}+(cosAsinB)^{2}\]

OpenStudy (anonymous):

the middle terms cancelled each other out when you added the two right sides together.

OpenStudy (anonymous):

Yes, that's where I got :) I just don't know how to carry on from there.

OpenStudy (anonymous):

Wait, I think you're missing one What of (sinAsinB)^2?

OpenStudy (anonymous):

group the firs and last terms and factor out cos^2A: (cosAcosB)^2 + (sinAsinB)^2 + (sinAcosB)^2 + (cosAsinB)^2 yes... you're right. almost done though.... are you up to this point? (double check for me)

OpenStudy (anonymous):

Yes, I'm up to this point. I double checked, and you only left out the (sinAsinB)^2. Thanks!! I hope you can get to the 1 and show me :)

OpenStudy (anonymous):

piece of cake from here....

OpenStudy (anonymous):

cosAcosB)^2 + (sinAsinB)^2 + (sinAcosB)^2 + (cosAsinB)^2 = [cosAcosB)^2 + (cosAsinB)^2] + [(sinAsinB)^2 + (sinAcosB)^2] = [cos^2A*cos^2B + cos^2A*sin^2B] + [sin^2A*sin^2B + sin^2A*cos^2B] = cos^2A*[cos^2B + sin^2B] + sin^2A*[sin^2B + cos^2B] = cos^2A*[1] + sin^2A*[1] = cos^2A + sin^2A = 1 . wheWWWWWWW

OpenStudy (anonymous):

i guess i gotta review my trig formulas!!!

OpenStudy (anonymous):

Thanks for all your help!

OpenStudy (anonymous):

nw

OpenStudy (anonymous):

Wait, how did you get rid of cos&2B + sin^2B?